OpenStudy (anonymous):

help please >.< (20x^2 - 13x + 2) ÷(4x - 1)

OpenStudy (anonymous):

\[4x-1\sqrt{(20x^2 - 13x + 2)}\]

OpenStudy (anonymous):

If you start with (20x^2 - 13x + 2) ÷(4x - 1) The first step you will want to factor our the numerator, so lets look at: (20x^2 - 13x + 2) we can factor this to: (4x-1)(5x-2) Now lets put this back in the original equation (4x-1)(5x-2)÷(4x - 1) so let us say (4x-1)=Y If we have Y(5x-2)/Y we can simplify this to get our final answer

OpenStudy (anonymous):

I am not sure what you did there @telijahmed but I do not think it is correct. If they are looking to simplify.

OpenStudy (anonymous):

still the same as you @edlevin just used the kindergarden method

OpenStudy (anonymous):

So.... what do i do after that?

OpenStudy (anonymous):

If you cancel the Y as I showed above, you should be left with (5x-2) as your final answer, did you want to solve for X also or just simplify it?

OpenStudy (anonymous):

OH. Okay, i get it. thank you :)

OpenStudy (anonymous):

\[4x-1\sqrt{(20x^2 - 13x + 2)}=5x-2\]

jimthompson5910 (jim_thompson5910):

you can also use polynomial long division to get what you see attached you'll get the same answer of 5x - 2

OpenStudy (anonymous):

Thank you guys ! :D